# Effective synchronization amid noise-induced chaos

### v1

@@ Abstract
Two remote agents with synchronized clocks may use them to act in concert and communicate. This necessitates some means of creating and maintaining synchrony. One method, not requiring any direct interaction between the agents, is exposing them to a common, environmental, stochastic forcing. This “noise-induced synchronization” only occurs under sufficiently mild forcing; stronger forcing disrupts synchronization. We investigate the regime of strong noise, where the clocks’ phases evolve chaotically. Using a simple realization of disruptive noise, we demonstrate effective synchronization. Namely, although the phases of the two clocks vary erratically, their probability distributions at a given moment become identical. We exploit this statistical synchrony to define an effective phase for an agent that closely agrees with the other’s phase. We discuss how this synchronization of ensembles broadens the range in which noise-induced synchronization can be used, and how it might be exploited in living systems.
Significance Statement
In ecology and neurobiology, independent living agents may anticipate one another’s behavior by means of a shared measure of time, e.g. simple chemical clocks requiring constant synchronization. We consider the effect of random events in the agents’ common environment that disrupt synchronization. We show that despite the resultant chaotic dynamics of their clocks, the two agents can independently analyze the recent history of the environmental noise to recover effective synchronization. This scheme offers a route to explain how coordinated behavior can emerge amid shared disruptive noise. An example is the puzzling large information transfer encoded in spike timings of sensory neurons.
@@ I Introduction
Living systems and man-made machines often depend on the cooperation amongst independent agents. This cooperation is often attained using chemical [1, 2, 3], electrical [4, 5, 6, 7], or mechanical [8] periodic processes (“clocks”), operating in each agent. Agents with synchronized clocks can exhibit emergent collective phenomena and are able to communicate [9, 10, 11, 12]. Since independent clocks inevitably lose synchronization over time, some means are needed in order to achieve and maintain synchronization.
@@
One such means is noise-induced synchronization, occurring when identical noninteracting oscillators are subjected to a common stochastic perturbation from their environment [10, 13, 14, 15, 16, 17, 18, 19, 20, 21]. It is distinct from synchronization due to interactions between the clocks and from phase locking unto an external clock. Instead, synchronization is enabled via the clocks’ shared ambient noise. As such, noise-induced synchronization is the only means of maintaining synchronization between noninteracting clocks in a generic environment. When the noise is sufficiently mild, it can bring and keep the two oscillators in synchrony over time. This has been well characterized experimentally in a range of physical contexts, such as electric circuits [18], quantum optics signaling [22], and even gene regulatory networks [23]. Synchronization can enable time-based communication between a sender and a receiver. If the sender emits an impulse when its clock is at a particular phase, the receiver can infer this phase by noting its own clock’s phase at the moment of the impulse.
@@
There is no guarantee that two oscillators will synchronize under a given noise. In general, synchronization requires the strength of the noise to lie below a threshold. Beyond this threshold, the relative phase of the two oscillators devolves into chaos [15, 21]. In the latter case, synchronization is lost and with it the means of time-based communication outlined above. Nevertheless, we show below that a form of synchronization persists in the statistical properties of these erratically evolving phases. It permits each agent to define an effective phase that agrees with that of other agents to a quantitatively-predictable extent, thus enabling effective synchronization.
@@
We demonstrate this behavior using a simplified type of noise, consisting of randomly timed impulses. For example, this may represent an electric shock applied to a collection of firing neurons [7, 4, 6] or a chemical affecting the cell cycle added into a cell culture [3], at stochastic times. We study the effect of these perturbations on a distribution of phases. Successive impulses create a sequence of phase distributions that in general depend on the initial distribution and the timings of the impulses. In below-threshold noise-induced synchronization, these distributions become narrow and unimodal, concentrated at a single phase. For noises above the threshold, the distributions remain spread over the entire phase circle [21].
@@
The above-threshold effective synchronization we report here arises by virtue of two properties: First, though the distributions change markedly with each impulse, we find that they become independent of the initial state. This means that two initially different oscillator ensembles become statistically identical when compared at a given instant. Second, the distributions can be tightly bunched into a few sharp peaks. Indeed, when the threshold of synchronization is approached from above, the distributions show a high degree of order, since their average entropy becomes arbitrarily negative [21].
@@
To demonstrate this effect, we follow two different initial phase distributions for two agents ${\mathrm{A}}$ and ${\mathrm{B}}$. We numerically determine their phase distributions $q^{(k)}_{{\mathrm{A}}}(\varphi)$ and $q^{(k)}_{{\mathrm{B}}}(\varphi)$ achieved after a sequence of $k$ impulses (“kicks”), which progressively converge with more kicks into a single distribution determined by the recent history of random waiting times. We propose an effective phase $\varphi_{\mathrm{f}}[q(\varphi)]$ for a distribution $q(\varphi)$ that each agent may estimate independently. We find that the effective phases $\varphi_{\mathrm{f}}$ of each agent ${\mathrm{A}}$ and ${\mathrm{B}}$ closely agree.
@@ II.1 Impulsive perturbations as phase maps
The effect of a kick on the clock can be compactly expressed via a phase-reduction prescription [19]: The oscillator traces a periodic loop in its dynamical manifold with a period $T$. We may label the points on this loop by a “phase position” $\varphi$ ranging from 00 to $1$, with a designated point assigned as the phase origin $\varphi\equiv 0$. Other points are assigned phase values equal to the time $\Delta t$ required for the oscillator to reach them from the phase origin: $\varphi\equiv\Delta t/T$. This construction is equivalent to finding the action-angle variables in Hamiltonian mechanics [24]. Thus, the phase of an oscillator that started at $\varphi^{(0)}$ advances at a constant speed, $\varphi^{(0)}+t/T$, where mod $\mathrm{mod}\,1$ is implied in any algebraic operation over the phases.
Every kick to an oscillator creates a fixed long-term shift in its time-dependent phase compared to the state of the oscillator if there had been no kick. This shift depends on the oscillator’s phase position $\varphi^{\prime}$ at the moment of the kick. All effects described below result from the rearrangement of the oscillator positions due to this position-dependent shift. For impulsive kicks, it suffices to know the shift of an oscillator kicked at $\varphi^{\prime}$, relative to the phase of an unkicked oscillator that was at the phase origin at the moment of the kick. This shifted phase, denoted $\psi$, is illustrated in Fig. 1. Since all phase points move around the orbit at the same constant speed, the phase difference $\psi$ for a given oscillator remains constant until the next kick. Evidently, if the kick has a negligible effect, the phase of an oscillator — initially equal to $\varphi$ — remains so, and thus $\psi(\varphi)=\varphi$. Otherwise, the “phase map” function $\psi(\varphi)$ encodes the complete information about the effect of the perturbation.
@@
We study the case where this deterministic perturbation $\psi(\varphi)$ is applied at random times. [^1] We assume that the waiting times in between kicks are sufficiently long so the oscillators have relaxed back to the limit cycle. For the case of possibly insufficiently-long times, see Ref. [29]. For simplicity, we may capture the effect of a $k$’th kick by observing the oscillators at an integer number of cycles after the latest kick. The unkicked position of the phase origin has continued its periodic motion and thus returns to the phase origin at such a moment. Likewise, the same phase is observed for the kicked oscillator after each completed cycle; we denote this phase as $\varphi^{(k)}$. Thus, the phase-map shift $\psi(\varphi^{\prime})$ from the unkicked phase origin is simply $\varphi^{(k)}$. To complete the iteration, we must express the phase position at the moment of the $k$th kick $\varphi^{\prime}$ in terms of the previous kick’s phase $\varphi^{(k-1)}$. During the random waiting time between the $k-1$’st and the $k$’th kick, the oscillator has gained a phase that we denote by $\beta^{(k)}$. Thus, $\varphi^{\prime}=\varphi^{(k-1)}+\beta^{(k)}$. Combining, we infer the effect of a single iteration,
@@
Due to the mod $\mathrm{mod}\,1$ constraint, only the fractional part of the waiting time in between kicks matters, which we denoted $\beta^{(k)}$. For simplicity, we assume that $\{\beta^{(1)},\beta^{(2)},\ldots\}$ are identical independent uniformly-distributed random phases, $\beta^{(k)}\in[0,1)$. Eq. (1), therefore, amounts to a Markov dynamics in the discrete “time” $k$, controlled by the phase map $\psi(\varphi)$ and the drawn $\beta^{(k)}$’s. Below, we compare the fates of two ensembles of oscillators exposed to the same sequence of $\beta^{(k)}$’s. [^2] We stress that all synchronization phenomena discussed in our work do not originate from phase distributions $q^{k}(\varphi)$ generated by an ensemble of noise histories $\{\beta^{(k)}\}$’s, but rather from the deterministic dynamics arising from a fixed history $\{\beta^{(k)}\}$ for an ensemble of initial states.
@@
where the “gain parameter” $A$ determines the strength of the perturbations; when $A=0$, $\psi(\varphi)=\varphi$, so the kick has no effect. We plot this $\psi(\varphi)$ for $A=9.32$ in SI Appendix, Fig. S4, solid red curve. For completeness, we have repeated all the paper’s quantitative analyses also for the quintic phase map, $\psi(\varphi)=\varphi+A^{\prime}\varphi(1-\varphi)(\Psi_{1}-\varphi)(\Psi_{2}-\varphi)(C-\varphi)$, where $C=(1-\Psi_{1})(1-\Psi_{2})/[\Psi_{1}\Psi_{2}+(1-\Psi_{1})(1-\Psi_{2})]$ is such that $(d\psi/d\varphi)|_{\varphi=1}=(d\psi/d\varphi)|_{\varphi=0}$. We will show the results for the cubic phase maps; we observed that the same conclusions apply to the quintic ones.
@@ II.2 Lyapunov exponent and synchronization
[…]
@@
Phase maps with $\Lambda<0$ tend on average to condense the phases. This means that oscillators that were subjected to the external force at common random waiting times $\beta^{(k)}$ are guaranteed to synchronize [17, 21]. On the other hand, $\Lambda>0$ suggests that neighboring oscillators tend to spread apart in phase within successive kicks. This means that oscillators subjected to such forcings are guaranteed to not asymptote to a synchronized state in general. Increasing the gain parameter $A$ ultimately causes $\Lambda$ to cross the threshold from negative to positive. This threshold distinguishes “mild” from “strong” perturbations.
@@ II.3 Probability distributions
Prior to the perturbation, the phase distribution remains unchanged up to rotation: $q^{(0)}(\varphi-t/T)$ (see illustration in SI Appendix, Fig. S1). We apply the same forcing (phase map $\psi$) after waiting times $\{\beta^{(k)}\}$ common to all oscillators. After the $k$’th kick, the ensemble will adopt the distribution $q^{(k)}(\varphi)$ according to
[…]
For $\Lambda<0$, synchronization manifests in a very narrow, unimodal (“single-peaked”) $q^{(k)}(\varphi)$’s which will only get narrower on average. In contrast, for $\Lambda>0$ one finds that $q^{(k)}(\varphi)$ continues to change erratically with each kick, with no ultimate convergence to narrow distributions. In Fig. 2, we show the typical evolution of two distributions under $\Lambda<0$ and $\Lambda>0$.
@@
Figure 2: A sequence of two, initially-different phase distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ (red) and $q^{(k)}_{\mathrm{B}}(\varphi)$ (blue) subjected to common noise in a typical realization of Eq. (1) and Eq. (2) with either $\Lambda=-0.141$ ($A=7.32$) or $\Lambda=0.141$ ($A=9.32$), as indicated. The initial distributions are uniform, $\varphi\sim\mathrm{U}[0.0,0.05)$ and $\varphi\sim\mathrm{U}[0.5,0.55)$, respectively. The distributions are shown on a circle so the periodicity mod $\,\mathrm{mod}\,1$ of the phase circle is apparent. The radial axis shows the distributions’ values $q(\varphi)$ on a log scale, where $q(\varphi)=1$ for the inner full circle and as indicated for the outer dotted circles. The kick numbers $k$ are shown inside the inner circle. The $\Lambda<0$ dynamics are synchronizing, so their distributions become centered around the same phase value and their width decreases over time on average. For $\Lambda>0$, the distributions evolve erratically. However, remarkably, the red and blue points co-locate; they converge unto the same sequence of sharply-multimodal distributions under common noise — this is the central finding of our work. The distributions are sampled by tracking $N=500$ initial phases via the nearest-neighbor distances; see Eq. (8). The entropies of the two distributions and the KLDs among them are shown in Fig. 3 for $\Lambda=0.141$ and SI Appendix, Fig. S6 for $\Lambda=-0.141$. A detailed view of the blue distribution at $k=100$ for $\Lambda=0.141$ is shown in SI Appendix, Fig. S7.
@@
In Sec. III, we show that after sufficiently many kicks, the two distributions become equal. This means that the same information is shared among the two agents, as both oscillators eventually follow the same statistics. To characterize this equivalence, we use the Kullback-Leibler divergence (KLD) [30] — a natural means of quantifying the divergence of the distribution $q_{\mathrm{A}}(\varphi)$ from $q_{\mathrm{B}}(\varphi)$,
[…]
@@
The convolution of Eq. (4) cannot be solved analytically. Instead, we implement Eq. (4) numerically by representing each $q^{(k)}(\varphi)$ as a sampling of $N$ discrete phases $\{\varphi_{1}^{(k)},\ldots,\varphi_{N}^{(k)}\}$, which by construction follow Eq. (1). This set of phases is used to estimate the entropies [31] and KLDs [32]. The details of the numerical estimators can be found in Sec. A.1. The numerical errors in their estimations for the samples sizes considered here are $\Delta S\sim 0.2$ and $\Delta D\sim 0.05$.

### v2

@@ Abstract
Two remote agents with synchronized clocks may use them to act in concert and communicate. This necessitates some means of creating and maintaining synchrony. One method, not requiring any direct interaction between the agents, is to expose them to a common, environmental, stochastic forcing. This “noise-induced synchronization” only occurs under sufficiently mild forcing; stronger forcing disrupts synchronization. We investigate the regime of strong noise, where the clocks’ relative phases evolve chaotically. Using a simple realization of disruptive noise, we demonstrate effective synchronization. First, although the relative phases of the two clocks varied erratically, we confirm that they became statistically independent of initial conditions and hence equivalent after a well-defined timescale. Second, we show that an agent can estimate an effective phase that closely agrees with the other’s phase. Thus, synchronization is practically attainable beyond the regime of conventional noise-induced synchronization. We finally discuss how it might be used in living systems.
Significance
In ecology and neurobiology, independent living agents may anticipate one another’s behavior by means of a shared measure of time, e.g., simple chemical clocks requiring constant synchronization. We consider the effect of random events in the agents’ common environment that disrupt synchronization. We show that despite the resultant chaotic dynamics of their clocks, the two agents can independently analyze the impact of the recent environmental noise on their clock phases to recover effective synchronization. This scheme offers a route to explain how coordinated behavior can emerge amid shared disruptive noise. An example is the puzzling large information transfer encoded in spike timings of sensory neurons.
@@ I Introduction
Living systems and man-made machines often depend on the cooperation amongst independent agents. This cooperation is often attained using chemical [1, 2, 3], electrical [4, 5, 6, 7, 8, 9], or mechanical [10] periodic processes (“clocks”), operating in each agent. Agents with synchronized clocks can exhibit emergent collective phenomena and are able to communicate [11, 12, 13, 14]. Since independent clocks inevitably lose synchronization over time, some means are needed in order to achieve and maintain synchronization.
@@
One such means is noise-induced synchronization, occurring when identical noninteracting oscillators are subjected to a common, nonlinear, stochastic perturbation from their environment [12, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. It is distinct from synchronization due to interactions between the clocks and from phase locking unto an external clock. Instead, synchronization is enabled via the clocks’ shared ambient noise. As such, noise-induced synchronization is the only means of maintaining synchronization between noninteracting clocks in a generic environment. When the noise is sufficiently mild, it can bring the two oscillators into synchrony over time. This has been well characterized experimentally in a range of physical contexts, such as electric circuits [21], quantum optics signaling [25], and even gene regulatory networks [26]. Synchronization can enable time-based communication between a sender and a receiver. If the sender emits an impulse when its clock is at a particular phase, the receiver can infer this phase by noting its own clock’s phase at the moment of the impulse.
@@
There is no guarantee that two oscillators will synchronize under a given noise. In general, synchronization requires the strength of the noise to lie below a threshold. Beyond this threshold, the relative phase of the two oscillators devolves into chaos [18, 24]. In the latter case, synchronization is lost and with it the means of time-based communication outlined above. Nevertheless, we show below that a form of synchronization persists in the statistical properties of these erratically evolving phases. It permits each agent to define an effective phase that agrees with that of other agents to a quantitatively-predictable extent, thus enabling them to behave as though synchronized.
@@
We demonstrate this behavior using a simplified type of noise, consisting of randomly timed impulses. For example, this may represent stochastically timed electric shocks delivered to a collection of firing neurons [7, 4, 6] or injections of a chemical affecting the cell cycle into a cell culture [3]. We study the effect of these perturbations on a distribution of phases. Successive impulses create a sequence of phase distributions that in general depend on the initial distribution and the timings of the impulses. In below-threshold noise-induced synchronization, these distributions become narrow and unimodal, concentrated at a single phase. For noises above the threshold, the distributions remain spread over the entire phase circle [24].
@@
The above-threshold effective synchronization we report here arises by virtue of two properties: First, though the distributions change markedly with each impulse, we find that they become independent of the initial state. This means that two initially different oscillator ensembles become statistically equivalent when compared at a given instant. Second, the distributions can be tightly bunched into a few sharp peaks. Indeed, when the threshold of synchronization is approached from above, the distributions show a high degree of order, since their average entropy becomes arbitrarily negative [24].
@@
To demonstrate this effect, we numerically follow two different initial phase distributions for two agents $\mathrm{A}$} and $\mathrm{B}$}. We find that the estimated Kullback-Leibler divergence among the two distributions, measuring their statistical difference, decreases with more kicks to zero after a characteristic kick number denoted $K_{\mathrm{m}}$. Thus, the phase distributions are determined from the most recent $K_{\mathrm{m}}$ kicks. We propose an effective phase $\varphi_{\mathrm{f}}$ for a given distribution that each agent may estimate independently. We find that the effective phases $\varphi_{\mathrm{f}}$ of each agent $\mathrm{A}$} and $\mathrm{B}$} closely agree.
@@ II.1 Impulsive perturbations as phase maps
The effect of a kick on the clock can be compactly expressed via a phase-reduction prescription [22]: The oscillator traces a periodic loop in its dynamical manifold with a period $T$. We may label the points on this loop by a “phase position” $\varphi$ ranging from 0 to $1$, with a designated point assigned as the phase origin $\varphi\equiv 0$. Other points are assigned phase values equal to the time $\Delta t$ required for the oscillator to reach them from the phase origin: $\varphi\equiv\Delta t/T$. This construction is equivalent to finding the action-angle variables in Hamiltonian mechanics [27]. Thus, the phase of an oscillator that started at $\varphi^{(0)}$ advances at a constant speed, $\varphi^{(0)}+t/T$, where $\mathrm{mod}\,1$ is implied in any algebraic operation over the phases.
Every kick to an oscillator creates a fixed long-term shift in its time-dependent phase compared to the state of the oscillator if there had been no kick. This shift depends on the oscillator’s phase position $\varphi^{\prime}$ at the moment of the kick. All effects described below result from the rearrangement of the oscillator positions due to this position-dependent shift. For impulsive kicks, it suffices to know the shift of an oscillator kicked at $\varphi^{\prime}$, relative to the phase of an unkicked oscillator that was at the phase origin at the moment of the kick. This shifted phase, denoted $\psi$, is illustrated in Fig. 1. Since all phase points move around the orbit at the same constant speed, the phase difference $\psi$ for a given oscillator remains constant until the next kick. Evidently, if the kick has a negligible effect, the phase of an oscillator — initially equal to $\varphi$ — remains so, and thus $\psi(\varphi)=\varphi$. Otherwise, the “phase map” function $\psi(\varphi)$ encodes the complete information about the long-term effect of the perturbation.
@@
We study the case where this deterministic perturbation $\psi(\varphi)$ is applied at random times. [^1] We assume that the waiting times in between kicks are sufficiently long so the oscillators have relaxed back to the limit cycle. For the case of possibly insufficiently-long times, see Ref. [32]. For simplicity, we may capture the effect of a $k$’th kick by observing the oscillators at an integer number of cycles after the latest kick. The unkicked position of the phase origin has continued its periodic motion and thus returns to the phase origin at such a moment. Likewise, the same phase is observed for the kicked oscillator after each completed cycle; we denote this phase as $\varphi^{(k)}$. Thus, the phase-map shift $\psi(\varphi^{\prime})$ from the unkicked phase origin is simply $\varphi^{(k)}$. To complete the iteration, we must express the phase position at the moment of the $k$th kick $\varphi^{\prime}$ in terms of the previous kick’s phase $\varphi^{(k-1)}$. During the random waiting time between the k1k-1’st and the $k$’th kick, the oscillator has gained a phase that we denote by $\beta^{(k)}$. Thus, $\varphi^{\prime}=\varphi^{(k-1)}+\beta^{(k)}$. Combining, we infer the effect of a single iteration,
@@
Due to the $\mathrm{mod}\,1$ constraint, only the fractional part of the waiting time in between kicks matters, which we denoted $\beta^{(k)}$. For simplicity, we assume that $\{\beta^{(1)},\beta^{(2)},\ldots\}$ are identical independent uniformly-distributed random phases, $\beta^{(k)}\in[0,1)$. Equation (1), therefore, amounts to a Markov dynamics in the discrete “time” $k$, controlled by the phase map $\psi(\varphi)$ and the drawn $\beta^{(k)}$’s. Below, we compare the fates of two ensembles of oscillators exposed to the same sequence of $\beta^{(k)}$’s. [^2] We consider synchronization of distinct initial phase distributions under the influence of a shared forcing, as opposed to identical initial phase distribution evolving under different forcings. We average over $\{\beta^{(k)}\}$’s only after computing differences between the two agents obtained with fixed $\{\beta^{(k)}\}$’s.
@@
where the “gain parameter” $A$ determines the strength of the perturbations; when $A=0$, $\psi(\varphi)=\varphi$, so the kick has no effect. We plot this $\psi(\varphi)$ for $A=9.32$ in SI Appendix, Fig. S4, solid red curve. For completeness, we have repeated the paper’s quantitative analyses also for the quintic phase map, $\psi(\varphi)=\varphi+A^{\prime}\varphi(1-\varphi)(\Psi_{1}-\varphi)(\Psi_{2}-\varphi)(C-\varphi)$, where $C=(1-\Psi_{1})(1-\Psi_{2})/[\Psi_{1}\Psi_{2}+(1-\Psi_{1})(1-\Psi_{2})]$ is such that $(d\psi/d\varphi)|_{\varphi=1}=(d\psi/d\varphi)|_{\varphi=0}$. We will show the results for the cubic phase maps; we observed that the same effective synchronization occurs for the quintic ones.
@@ II.2 Lyapunov exponent and synchronization
$\Lambda=\int_{0}^{1}d\varphi\ln\left|\frac{d\psi}{d\varphi}\right|.$ (3)
@@
Phase maps with $\Lambda<0$ tend to condense the phases. This means that oscillators that were subjected to the external force at common random waiting times $\beta^{(k)}$ are guaranteed to synchronize [20, 24]. On the other hand, $\Lambda>0$ suggests that neighboring oscillators tend to spread apart in phase over successive kicks. This means that oscillators subjected to such forcings are guaranteed to not asymptote to a synchronized state in general. Increasing the gain parameter $A$ ultimately causes $\Lambda$ to cross the threshold from negative to positive. This threshold distinguishes “mild” from “strong” perturbations.
@@ II.3 Probability distributions
Prior to the perturbation, the phase distribution remains unchanged up to rotation: $q^{(0)}(\varphi,t)=q^{(0)}(\varphi-t/T)$ (see illustration in SI Appendix, Fig. S1). We apply the same forcing (phase map $\psi$) after waiting times $\{\beta^{(k)}\}$ common to all oscillators. After the $k$’th kick, the ensemble will adopt the distribution $q^{(k)}(\varphi)$ according to
$q^{(k)}(\varphi)=\int_{0}^{1}d\varphi^{\prime}q^{(k-1)}(\varphi^{\prime})\delta(\varphi-\psi(\varphi^{\prime}+\beta^{(k)})).$ (4)
For $\Lambda<0$, synchronization manifests in a very narrow, unimodal (“single-peaked”) $q^{(k)}(\varphi)$’s which continually become narrower on average. In contrast, for $\Lambda>0$ one finds that $q^{(k)}(\varphi)$ continues to change erratically with each kick, with no ultimate convergence to narrow distributions. In Fig. 2, we show the typical evolution of two distributions under $\Lambda<0$ and $\Lambda>0$.
@@
Figure 2: A sequence of two, initially-different phase distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ (red) and $q^{(k)}_{\mathrm{B}}(\varphi)$ (blue) subjected to common noise in a typical realization of Eq. (1) and Eq. (2) with either $\Lambda=-0.141$ ($A=7.32$) or $\Lambda=0.141$ ($A=9.32$), as indicated. The initial distributions are uniform, $\varphi\in[0.0,0.05)$ and $\varphi\in[0.5,0.55)$, respectively. The distributions are shown on a circle so the periodicity $\mathrm{mod}\,1$ of the phase circle is apparent. The radial axis shows the distributions’ values $q(\varphi)$ on a log scale, where $q(\varphi)=1$ for the inner full circle and as indicated for the outer dotted circles. The kick numbers $k$ are shown inside the inner circle. The $\Lambda<0$ dynamics are synchronizing, so their distributions become centered around the same phase value and their width decreases over time on average. For $\Lambda>0$, the distributions evolve erratically. However, remarkably, the red and blue points co-locate; they converge unto the same set of sharply-multimodal distributions under common noise — this property enables the effective synchronization demonstrated below. The distributions are sampled by tracking $N=500$ initial phases via the nearest-neighbor distances; see Eq. (8). The entropies of the two distributions and the KLDs among them are shown in Fig. 3 for $\Lambda=0.141$ and SI Appendix, Fig. S6 for $\Lambda=-0.141$. A detailed view of the blue distribution at $k=100$ for $\Lambda=0.141$ is shown in SI Appendix, Fig. S7.
@@
In Sec. III, we show that after sufficiently many kicks, the two distributions become operationally equivalent. This means that the same information is shared among the two agents, as both oscillators eventually follow the same statistics. To characterize this equivalence, we use the Kullback-Leibler divergence (KLD) [33] — a natural means of quantifying the divergence of the distribution $q_{\mathrm{A}}(\varphi)$ from $q_{\mathrm{B}}(\varphi)$,
$D({\mathrm{A}}\|{\mathrm{B}})\equiv\int_{0}^{1}d\varphi q_{\mathrm{A}}(\varphi)\ln\frac{q_{\mathrm{A}}(\varphi)}{q_{\mathrm{B}}(\varphi)}.$ (5)
@@
The convolution of Eq. (4) poses numerous issues: First, it cannot be carried out analytically for an arbitrary $q^{(k-1)}(\varphi)$. Second, each extremum in $\psi(\varphi)$, denoted $\varphi=\varphi_{\mathrm{m}}$, imparts an integrable singularity on $q^{(k)}(\varphi)$ of the form $\sim|\varphi-\psi(\varphi_{\mathrm{m}})|^{-1/2}$; these keep accumulating with each kick. Further, determining $q(\varphi)$ in practice requires sampling it with a finite number $N$ of samples. Thus, we sample the chosen initial distribution $q^{(0)}(\varphi)$ by $N$ samples $\{\varphi^{(0)}_{1},\varphi^{(0)}_{2},...\varphi^{(0)}_{N}\}$ drawn from $q^{(0)}(\varphi)$. Each of these $N$ samples changes under a kick via the phase map, according to Eq. (1). By repeatedly applying Eq. (1), we obtain the $\{\varphi^{(k)}_{1},\varphi^{(k)}_{2},...\varphi^{(k)}_{N}\}$, corresponding to a sample of $q^{(k)}(\varphi)$. We may then use standard tools to estimate $S^{(k)}$ [34] and $D^{(k)}(A||B)$ [35] in terms of the $\{\varphi^{(k)}_{n}\}$, as described in Appendix A.1. (The numerical uncertainty owing to discrete sampling in our simulations was of order $0.2$ for $S$ and $0.05$ for $D$ with $N=500$.) As we show in this work, the distributions of Fig. 2, obtained by such a discrete sampling, carry sufficient information about a synchronized phase notwithstanding the above singularities.
III Convergent distributions